# A System Level Approach to LQR Control of the Diffusion Equation

### v1

@@ VI Equivalence to LQR Control
We demonstrate that the solution to our $H_{2}$ design problem indeed recovers the solution to the standard infinite-horizon LQR problem for the diffusion equation.
@@
Let $(\mathcal{F}\psi_{0})(\kappa)=1$ for all $\kappa$.  [^2] Note that the “response” (weak solution) of a system to this initial condition is equal to the Green’s Function of the linear PDE describing the system dynamics, and that $\mathcal{F}$ here is the extension of the Fourier transform to the space of distributions. Then the standard LQR problem (LABEL:eq:LQR) is equivalent to the $H_{2}$ problem of interest (LABEL:eq:H2_formulation).
@@
First, we compute the solution to the LQR formulation (LABEL:eq:LQR) using standard approaches [16, 14]. We solve the operator-valued Ricatti equation
@@
The unique solution $\hat{F}_{\kappa}>0$ to (46) is $\hat{F}_{\kappa}=-\alpha\kappa^{2}+\sqrt{\alpha^{2}\kappa^{4}+\gamma^{2}},$ and the optimal state feedback gain is thus given by
$K=\mathcal{F}M_{\hat{K}}\mathcal{F}^{-1},~~\hat{K}_{\kappa}=-\hat{F}_{\kappa}=\alpha\kappa^{2}-\sqrt{\alpha^{2}\kappa^{4}+\gamma^{2}}.$ (47)
@@
$\hat{K}_{\kappa}(s)=\frac{\hat{\Phi}^{u}_{\kappa}(s)}{\hat{\Phi}^{\psi}_{\kappa}(s)}=\alpha\kappa^{2}-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}.$ (48)
This is indeed equal to the static state feedback gain $\hat{K}$ obtained from the Riccati equation solution. Details on the equivalence of the cost of the two formulations (LABEL:eq:CL_formulation) and (LABEL:eq:LQR) are provided in Appendix -B. ∎
We know that $K$ will take the form of a spatial convolution operator
@@
We compare the static controller $K$ to the dynamic controller implementation (43) depicted in Fig. 1. As previously noted, the transfer function $\left(I-G(s){\Phi}^{\psi}(s)\right)$ is strictly proper, thus its corresponding convolution kernel has only a dynamic part, which we denote as $f_{\Phi^{\psi}}(t,\theta)$. The transfer function $G(s)\Phi^{u}(s)$ has a corresponding convolution kernel with both a static and dynamic part, which we denote as $h_{\Phi^{u}}(\theta),f_{\Phi^{u}}(t,\theta)$, respectively. The static (feedthrough) component of $G(s)\Phi^{u}(s)$ is obtained by evaluating its Fourier symbol in the limit as $s\to\infty$,
$\mathcal{F}^{-1}\lim_{s\to\infty}\hat{G}(s,\kappa)\hat{\Phi}^{u}(s,\kappa)=h_{\Phi^{u}}(\theta)=(\mathcal{F}^{-1}\hat{K})(\theta),$
which shows that the static part of the implementation (43) recovers the static LQR gain, as expected. The “dynamic components” of the systems $G(s)\Phi^{u}(s)$ and $\left(I-G(s){\Phi}^{\psi}(s)\right)$ cancel each other out, leaving only the static feedback policy $K.$ This is confirmed in the plot of $h_{\Phi^{u}}(\theta),f_{\Phi^{u}}(t,\theta),f_{\Phi^{\psi}}(t,\theta)$ shown in Fig. 2.
@@ VII Incorporation of Spatio-Temporal Decay Rate Constraints
Without constraints, the closed-loop formulation of the $H_{2}$ problem (LABEL:eq:CL_formulation) is equivalent to a standard LQR problem. The advantage of this closed-loop formulation is that it remains convex when spatio-temporal constraints are incorporated. A similar problem has been studied in [2], where necessary and sufficient conditions for convexity were dependent on “funnel causality” properties of the underlying system and desired constraint. Notably, our formulation does not rely on these conditions and thus admits a broader range of constraints while preserving convexity.
@@
$\inf_{\Phi^{\psi},\Phi^{u}\in H_{\infty}}~\left\|C\Phi^{\psi}+D\Phi^{u}\right\|_{H_{2}}$ (50)
@@
$\|\rho(\cdot,t)-\rho_{T}(\cdot,t)\|_{L_{2}(\mathbb{T})}<\epsilon(t)$
@@
and $r\leq\pi$ is the truncation radius that encodes communication limitations. This constraint ensures that the ‘tails’ of $\rho$ are sufficiently small, potentially enforcing spatial locality. Choosing $\epsilon(t)$ to grow in $t$ allows $\rho$ to be less localized in space as time increases. This set $\mathcal{S}$ is convex, and guarantees that a spatially truncated version of controller will perform well using small gain arguments.
@@ VIII Conclusion
It was shown that the infinite-horizon LQR problem for the diffusion equation over the unit circle can be reformulated as an $H_{2}$ problem over the system’s closed-loop responses, analogous to the finite-dimensional System Level Synthesis methodology. Implementation of the resulting controller using the closed-loop responses was shown to preserve stability of the closed-loop system. It was further illustrated that this novel formulation admits constraints on the closed-loop responses while preserving convexity. Generalizations of the presented results to a class of linear distributed parameter systems over both compact and unbounded spatial domains are the subject of ongoing work. Of particular interest is comparing analytical results from this framework to the computational approach of [9].

### v2

@@ VI Equivalence to LQR Control
We demonstrate that the solution to our closed-loop $\mathcal{H}_{2}$ design problem (LABEL:eq:CL_formulation) recovers the solution to the standard infinite-horizon LQR problem for the diffusion equation.
@@
The standard LQR problem (LABEL:eq:LQR) and the $\mathcal{H}_{2}$ problem of interest (LABEL:eq:CL_formulation) have the same solution $K^{\rm opt}$.
@@
First, we compute the solution to the LQR formulation (LABEL:eq:LQR) using standard approaches [2, 7]. We solve the operator-valued Riccati equation
@@
The unique solution $\hat{F}_{\kappa}>0$ to (38) is $\hat{F}_{\kappa}=-\alpha\kappa^{2}+\sqrt{\alpha^{2}\kappa^{4}+\gamma^{2}},$ and the optimal state feedback gain is thus given by $K^{\rm LQR}=\mathcal{F}M_{\hat{K}^{\rm LQR}}\mathcal{F}^{-1}$ where
$\hat{K}^{\rm LQR}_{\kappa}=-\hat{F}_{\kappa}=\alpha\kappa^{2}-\sqrt{\alpha^{2}\kappa^{4}+\gamma^{2}}.$ (39)
@@
$\hat{K}^{\mathcal{H}_{2}}_{\kappa}(s)=\frac{\hat{\Phi}^{u}_{\kappa}(s)}{\hat{\Phi}^{\psi}_{\kappa}(s)}=\alpha\kappa^{2}-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}.$ (40)
This is indeed equal to the static state feedback gain $\hat{K}^{LQR}$, thus $K^{\rm LQR}=K^{\mathcal{H}_{2}}=K^{\rm opt}.$
We know that $K^{\rm opt}$ will be a spatial convolution operator
@@
We compare the static controller $K^{\rm opt}$ to the dynamic controller implementation (35) depicted in Fig. 1. As previously noted, the transfer function $\left(I-G(s){\Phi}^{\psi}(s)\right)$ is strictly proper, thus its corresponding convolution kernel has only a dynamic part, which we denote as $f_{\Phi^{\psi}}(t,\theta)$. The transfer function $G(s)\Phi^{u}(s)$ has a corresponding convolution kernel with both a static and dynamic part, which we denote as $h_{\Phi^{u}}(\theta),f_{\Phi^{u}}(t,\theta)$, respectively. The static (direct feedthrough) component of $G(s)\Phi^{u}(s)$ is obtained by evaluating its Fourier symbol in the limit as $s\to\infty$,
$\mathcal{F}^{-1}\lim_{s\to\infty}\hat{G}(s,\kappa)\hat{\Phi}^{u}(s,\kappa)=h_{\Phi^{u}}(\theta)=(\mathcal{F}^{-1}\hat{K}^{\rm opt})(\theta),$
which shows that the static part of the implementation (35) recovers the static LQR gain, as expected. The “dynamic components” of the systems $G(s)\Phi^{u}(s)$ and $\left(I-G(s){\Phi}^{\psi}(s)\right)$ cancel each other out, leaving only the static feedback policy $K^{\rm opt}.$ This is confirmed in the plot of $h_{\Phi^{u}}(\theta),f_{\Phi^{u}}(t,\theta),f_{\Phi^{\psi}}(t,\theta)$ shown in Fig. 2.
@@ VII Incorporation of Spatio-Temporal Decay Rate Constraints
Without constraints, the closed-loop formulation of the $\mathcal{H}_{2}$ problem (LABEL:eq:CL_formulation) is equivalent to a standard LQR problem. The advantage of this closed-loop formulation is that it remains convex when spatio-temporal constraints are incorporated. A similar problem has been studied in [3], where necessary and sufficient conditions for convexity were dependent on “funnel causality” properties of the underlying system and desired constraint. Notably, our formulation does not rely on these conditions and thus admits a broader range of constraints while preserving convexity.
@@
$\operatorname*{arg\,min}_{\Phi^{\psi},\Phi^{u}\in H_{\infty}}~\left\|C\Phi^{\psi}+D\Phi^{u}\right\|_{\mathcal{H}_{2}}$ (42)
@@
$\|\rho(\cdot,t)-\rho_{T}(\cdot,t)\|_{L_{2}(\mathbb{T})}<\epsilon(t)$ (43)
@@
and $r\leq\pi$ is the truncation radius that encodes communication limitations. This constraint ensures that the ‘tails’ of $\rho$ are sufficiently small, potentially enforcing spatial locality. Choosing $\epsilon(t)$ to grow in $t$ allows $\rho$ to be less localized in space as time increases. While constraining $\rho$ to have compact support as in [5] admits desirable computational properties, the decay constraint (43) is perhaps more physically relevant. This set $\mathcal{S}$ is convex, and guarantees that a spatially truncated version of controller will perform well using small gain arguments.
@@ VIII Conclusion
It was shown that the infinite-horizon LQR problem for the diffusion equation over the unit circle can be reformulated as an $\mathcal{H}_{2}$ problem over the system’s closed-loop responses, analogous to the finite-dimensional System Level Synthesis methodology. An implementation of this optimal controller built from these closed-loop responses was shown to preserve internal stability. It was further illustrated that this closed-loop formulation admits a class of spatio-temporal constraints on the closed-loop responses while preserving convexity. Generalizations of this work to general linear distributed parameter systems are the subject of ongoing work. Of particular interest include incorporation of boundary conditions through an embedding approach [9] or alternate representations [17]; and comparison of analytical results from this work to the computational approach of [6].